# Planning in a recurrent neural network that plays Sokoban

### v2

@@ 6 Related work
Interpreting agents and planning.
Several works have attempted to find the mechanism by which a simple neural network does planning in mazes [Mini et al., 2023, Knutson et al., 2024, Brinkmann et al., 2024], gridworlds [Bloom and Colognese, 2023], and graph search [Ivanitskiy et al., 2023]. Men et al. [2024] investigates LLM reasoning in a simple block-stacking task. We believe the DRC we present is a clearer example of an agent than what these works focus on, and should be similarly possible to interpret.
Concurrent work [Bush et al., 2024] interprets another DRC Sokoban agent and can also predict future actions. That team invented grid-wise probes and the Agent-Directions and Box-Directions targets.
Other works interpret superhuman game-playing agents. Jenner et al. [2024] find evidence of lookahead in Leela Chess Zero [Team, 2018], without predicting future actions. McGrath et al. [2021], Schut et al. [2023] find interpretable concepts in AlphaZero [Silver et al., 2017].
Probing for world models.
Planning usually requires a model of the world. Li et al. [2023], Nanda et al. [2023b], Karvonen [2024] examine agents trained to predict possible moves in games, and are able to probe for the board state. Wijmans et al. [2023] also probe the state of navigation neural networks. Gurnee and Tegmark [2023] find a representation of positions in Earth for location-associated tokens in an LLM.
Goal misgeneralization and mesa-optimizers for alignment.
From the alignment perspective, AIs optimizing monomaniacally for a goal have long been a concern (e.g. Yudkowsky, 2006, Omohundro, 2008; see the preface of Russell, 2019). In a machine learning paradigm [Hubinger et al., 2019], the goal of the training system is not necessarily optimized; instead, the NN may optimize for a related or different goal [Di Langosco et al., 2022, Shah et al., 2022] or for no goal at all.
Chain-of-thought faithfulness.
Large language models use chain of thought, but are they faithful to it, or do they think about their future actions in other ways [Lanham et al., 2023, Pfau et al., 2024]? One could hope that LLMs perform all long-term reasoning in plain English, allowing unintended human consequences to be easily monitored, as in Scheurer et al. [2023].
Ethical treatment of AIs.
Do AIs deserve moral consideration? Schwitzgebel and Garza [2015] argue that very human-like AIs are conceivable and clearly deserve rights. Tomasik [2015] suggests that most AIs deserve at least a little consideration, like biological organisms of any species [Singer, 2004]. But what does it mean to treat an AI ethically? Daswani and Leike [2015] argue that the way to measure pleasure and pain in a reinforcement learner is not by its absolute amount of return, but rather by the temporal difference (TD) error: the difference between its expectations and the actual return it obtained. If the internals of the NN have a potentially different objective [Hubinger et al., 2019, Di Langosco et al., 2022], then the TD error should come from a place other than the critic head. Research into learned search algorithms is an early step toward finding the learned-reward internal TD error, if it exists. This could be a higher-assurance complement to simply asking the AI [Perez and Long, 2023].
Fully reverse engineering small networks.
Many works reverse engineer all of, or most of, a small NN that does an algorithmic task [Nanda et al., 2023a, Chughtai et al., 2023, Zhong et al., 2023, Quirke and Barez, 2023].
Neural network architectures that reason.
Many papers try to enhance NN thinking by altering the training setup or architecture. Schwarzschild et al. [2021b, a], Bansal et al. [2022] improve and evaluate a specific kind of RNN training that produces algorithmic thinking, though Knutson et al. [2024] argue that they do not generalize enough and thus do not implement the correct algorithm. Other works endow RNNs with variable amounts of computation per step, with [Chung et al., 2024] or without [Graves, 2016] an explicit world model.
@@ 7 Conclusion
We replicated a small RL-based Sokoban agent, whose performance benefits from increased compute in the form of externally imposed no-op steps [Guez et al., 2019]. We find that the agent learns to explicitly exploit this capability by pacing, which substitutes for the external no-op thinking steps. Extra computation helps with longer levels and those that require non-myopic thinking.
We show that the agent causally represents plans (sequences of actions to take). The most fundamental plan representation we find is the box-direction probes, which can predict and control where the agent pushes a box. By building up a chain of box probe steps over time, the agent is able to figure out how to place boxes to solve the level. Secondarily, we find agent-direction probes that take lower precedence than the box-direction probes, but can control the agent moves in isolated setups. We additionally find that the box probe chains get longer and longer with more thinking steps (Figure 5).
Finally, we find that extracting a decision from the plan takes place almost entirely within the ConvLSTM, with the MLP doing little more than decoding it. To demonstrate this, we show that the agent can generalize to levels bigger than what it had ever seen during training by manually replacing its MLP decoder based on our mechanistic understanding.
We hope the model organism we trained and open-source will catalyze research into NNs that learn to plan, and to understand their inner objectives.
Acknowledgements
We would like to thank ChengCheng Tan and Siao Si Loo for editing this paper, and the rest of the FAR team for general support. We also thank Alex F. Spies and anonymous reviewers for improving the paper with their comments and edits. We are also very grateful for the many questions about their RL setup that Arthur Guez and Timothy Lillicrap answered, as well as inspiration from Stephen Chung and Scott Goodfriend.
Reproducibility Statement
We provide the code and the trained models in the supplementary materials. We open-source our code and trained model(s). Additionally, sections 2, 4.1 and appendices A, D, and E describe the specific experimental details and settings we used to carry out our experiments.
@@ A.1 Training hyperparameters
All networks were trained with the same hyperparameters, which were tuned on a combination of the ResNet and the $\text{DRC}(3,3)$. These are almost exactly the same as Guez et al. [2019], allowing for taking the mean of the per-step loss instead of the sum.
@@ A.3 Learning curve comparison
It is difficult to fully replicate the results by Guez et al. [2019]. Chung et al. [2024] propose an improved method for RL in planning-heavy domains. They employ the IMPALA $\text{DRC}(3,3)$ as a baseline and plot its performance in Chung et al. [2024, Figure 5]. They plot two separate curves for $\text{DRC}(3,3)$: that from Guez et al. [2019], and a decent replicated baseline. The baseline is considerably slower to learn and peaks at lower performance.
We did not innovate in RL, so were able to spend more time on the replication. We compare our replication to Guez et al. [2019] in Figure 7, which shows that the learning curves for $\text{DRC}(3,3)$ and ResNet are compatible, but not the one for DRC(1,1). Our implementation also appears much less stable, with large error bars and large oscillations over time. We leave addressing that to future work.
The success rate in Figure 7 is computed over 1024 random levels, unlike the main body of the paper. Table 5 reports test and validation performance for the DRC and ResNet seeds which we picked for the paper body.
@@
Figure 7: Success rate for Test-unfiltered and Validation-medium levels vs. environment steps of training. Each architecture has 5 random seeds, the solid line is the pointwise median and the shaded area spans from the minimum to the maximum. The dotted lines are data for the performance of architectures extracted from the [Guez et al., 2019] PDF file. The values are slightly different from what Figures 5 and 3 report because they are calculated on a random sample of 1024 levels (24 levels are repeated for test-unfiltered, which only has 1000 levels).
@@
Appendix B Generalizing the $\text{DRC}(3,3)$ to larger levels
@@
We license the levels by Þorsteinsson [2009] as GPLv3, and make them available in the supplementary material. The performance of the $\text{DRC}(3,3)$ on each set of levels is in Table 6. The sets with higher scores are also those which humans find easier: for example, “Dimitri & Yorick” was made for children by Jacques Duthen and consists entirely of relatively small levels (largest is $12\times 10$) with at most 5 boxes. The sets where the $\text{DRC}(3,3)$ solves nothing are also difficult for the authors of this paper.
@@
We tried giving the $\text{DRC}(3,3)$ from 2 to 128 extra thinking steps for these levels, incrementing in powers of two. For most sets, we find some benefit to 2-4 thinking steps, but no more. The sole exception is XSokoban, which contains one level (Figure 2, right) which requires 128 steps of thinking to solve.
@@
Table 6: Performance of the $\text{DRC}(3,3)$ on each set of levels by [Þorsteinsson, 2009]. The “Max solved” columns represent the proportion of levels solved at the number of steps in the “at steps” column, which is the highest solved proportion for each number of thinking steps tried.
All levels Levels larger than $10\times 10$
Level collection # solved Max slv. max at # solved Max slv. max at
Dimitri & Yorick 61 83.6% 88.5% 16 0
Sokoban Jr. 1 60 80.0% 81.7% 2 19 73.7% 73.7% 0
Howard’s 3rd set 40 60.0% 62.5% 4 1 0.0% 0.0% 0
Simple sokoban 61 54.1% 60.7% 16 51 47.1% 54.9% 16
Sokoban Jr. 2 54 48.1% 50.0% 2 40 45.0% 47.5% 2
Microban 155 23.9% 24.5% 4 17 5.9% 11.8% 2
Deluxe 55 21.8% 23.6% 2 1 0.0% 0.0% 0
Sokogen 990602 78 20.5% 23.1% 8 0
Yoshio Automatic 52 13.5% 17.3% 2 0
Sasquatch III 16 6.2% 6.2% 0 8 0.0% 0.0% 0
Howard’s 1st set 100 5.0% 6.0% 2 54 0.0% 0.0% 0
Howard’s 2nd set 40 5.0% 10.0% 4 22 0.0% 0.0% 0
Microcosmos 40 5.0% 7.5% 4 0
Still more levels 35 2.9% 2.9% 0 34 2.9% 2.9% 0
Sasquatch IV 36 2.8% 2.8% 0 20 0.0% 0.0% 0
Xsokoban 40 2.5% 5.0% 128 39 2.6% 5.1% 128
Sasquatch 49 2.0% 4.1% 4 39 0.0% 2.6% 4
@@
Appendix C Case Studies

### v3

@@ 7 Related Work
Our work contributes to the growing field of mechanistic interpretability focused on understanding reasoning and planning processes within neural networks. While the internal mechanisms of complex agents remain largely opaque, several lines of research provide context for our investigation.
Interpreting planning and reasoning in neural networks.
Prior studies have investigated planning mechanisms in simpler settings like mazes [Mini et al., 2023, Knutson et al., 2024, Brinkmann et al., 2024], gridworlds [Bloom and Colognese, 2023], and graph search problems [Ivanitskiy et al., 2023]. Others have explored reasoning in Large Language Models (LLMs) on tasks such as block-stacking [Men et al., 2024]. These works often focus on identifying representations of state or specific concepts, rather than studying the representation involved in planning. Researchers have also investigated sophisticated game-playing agents like Leela Chess Zero [Jenner et al., 2024] and AlphaZero [McGrath et al., 2021, Schut et al., 2023], uncovering evidence for lookahead mechanisms or high-level local concepts important for the game. Our work builds on these efforts but specifically targets the explicit representation and causal role of plans (complete sequences of future actions, Definition 1.1) within the DRC network trained on Sokoban, a game demanding complex, foresightful planning.
Sokoban planning and the DRC agent.
The DRC network trained on Sokoban provides a compelling case study of long-term deliberate planning owing to its ability to utilize extra thinking time [Guez et al., 2019]. Our work aims to understand and explain this behavior by interpreting the DRC’s internal states. Concurrently with our research, Bush et al. [2025] also interpreted a DRC agent trained on Sokoban, by introducing probes similar to ours that predict future box and agent movements and found evidence for plan representations (similar to our finding 1). Our work complements and extends theirs in several key ways. Bush et al. [2025] performed causal interventions on $200$ handcrafted toy levels, whereas we perform more rigorous causal tests on $24k$ datapoints sampled from difficult standard benchmark levels and devise an intervention methodology that edits the representations more precisely (Section 3). Additionally, our work yields novel findings: finding 4 offers further evidence from probe-guided OOD generalization that DRC’s planning capability generalizes to larger levels not supported by the architecture, while finding 3 on pacing behavior provides explanation for the emergence of the performance improvements through “thinking time”.
Probing and intervention methods.
Our methodology relies on probing internal activations to decode information embedded linearly and intervening on these activations to test causality. Linear probing has been widely successful in finding representations of game states [Li et al., 2023, Nanda et al., 2023b, Karvonen, 2024] or spatial information [Wijmans et al., 2023] in various networks trained on game sequences.
Adaptive computation for reasoning.
The finding that the DRC’s plan improves with computation (Finding 2) and that the agent learns to "pace" (Finding 3) connects to research on improving neural network reasoning by altering training setup or architecture and leveraging adaptive computation time in RNNs [Schwarzschild et al., 2021b, a, Bansal et al., 2022]. [Graves, 2016] and models that explicitly vary computation [Chung et al., 2024]. However, Knutson et al. [2024] argue that they lack generalization and fail to implement correct algorithms. Other works explore models that adjust computation per step, either with an explicit world model [Chung et al., 2024] or without one [Graves, 2016]. Our work suggests such adaptive strategies for long-term planning and reasoning can be learned implicitly using model-free RL, mirroring the recent advancements of LLMs leveraging test-time compute for improved reasoning [DeepSeek-AI et al., 2025, OpenAI, 2024].
Goal misgeneralization and mesa-optimization.
Understanding how neural networks perform tasks requiring sequential reasoning is crucial for ensuring transparency and safety of AI systems. In AI alignment, goal misgeneralization occurs when a neural network optimizes for unintended proxy objectives instead of its intended goal (e.g. Yudkowsky, 2006, Omohundro, 2008; see Russell, 2019). This risk relates to mesa-optimization, where a network develops internal objectives that diverge from training targets [Hubinger et al., 2019]. Recent work has shown that networks may pursue goals misaligned—or even directly opposed—to those intended by their designers [Di Langosco et al., 2022, Shah et al., 2022]. Mechanistic interpretability of AI agents is a crucial tool for understanding how neural networks internally represent goals and plans and for identifying and fixing potential misalignments.
We discuss more related work in Appendix I.
@@ 8 Conclusion
Building on prior and concurrent work on search-like behaviors of DRC agents [Guez et al., 2019] and plan representations [Bush et al., 2025], we provide detailed evidence that DRC agents represent causal plans through spatial encodings of future box movements and refine them over time.
We describe how extra thinking steps increase success rate, by preventing greedy actions before the plan stabilizes into a longer-term solution. We find the agent takes advantage of this by “pacing” on-distribution to get enough time to stabilize the plan. By substituting the network output module with much simpler probes, we demonstrate how a mechanistic understanding can enable generalization beyond the agent’s training distribution.
These findings advance the understanding of planning dynamics in neural networks and lay the groundwork for more transparent, safer and interpretable AI systems.
Impact Statement
This research into interpretability can make models more transparent, which helps in making models predictable, easier to debug and ensure they conform to specifications.
Specifically, we train and open-source a model organism which is planning, and analyze it somewhat; we hope this will catalyze further research on identifying, evaluating and understanding what goal a model has. We hope that directly identifying a model’s goal lets us monitor for and correct goal misgeneralization [Di Langosco et al., 2022].
Goal evaluation also has implications for AI welfare. Knowing the goals of a model would also make it possible to know whether these goals are being satisfied enough, and thus perhaps give a way to evaluate whether or not a model is happy (Appendix I), assuming it is a moral patient.
@@ A.1 Training hyperparameters
All networks were trained with the same hyperparameters, tuned on both ResNet and the $\text{DRC}(3,3)$. These largely match Guez et al. [2019], except we take the mean per-step losses instead of summing.
@@ A.3 Learning curve comparison
Replicating the results of Guez et al. [2019] proved challenging. Chung et al. [2024] propose an improved method for RL in planning-heavy domains. They employ the IMPALA $\text{DRC}(3,3)$ as a baseline and plot its performance in Chung et al. [2024, Figure 5]. They plot two separate curves for $\text{DRC}(3,3)$: that from Guez et al. [2019], and a decent replicated baseline. The baseline is considerably slower to learn and peaks at lower performance.
We did not innovate in RL, so were able to spend more time on the replication. We compare our replication to Guez et al. [2019] in Figure 8, which shows that the learning curves for $\text{DRC}(3,3)$ and ResNet are compatible, but not the one for DRC(1,1). Our implementation exhibits reduced stablility, with large error bars and pronounced oscillations over time. We defer addressing this to future work. Table 3 reports test and validation performance for the DRC and ResNet seeds which we picked for the paper body.
@@
Figure 7: Success rate of $\text{DRC}(3,3)$ on Validation-medium levels across forced thinking steps at the start of the episode. Increasing thinking steps increases performance.
Figure 8: Success rate for Test-unfiltered and Validation-medium levels vs. environment steps of training. Each architecture has 5 random seeds, the solid line is the pointwise median and the shaded area spans from the minimum to the maximum. The dotted lines are data for the performance of architectures extracted from the [Guez et al., 2019] PDF file.
@@
Appendix B Results on $\text{DRC}(1,1)$ and ResNet
B.1 $\text{DRC}(1,1)$
In this section, we show that the $\text{DRC}(1,1)$ network, which only has a single recurrent convolutional block that runs a single forward pass per environment step, also exhibits similar planning behavior to the $\text{DRC}(3,3)$ network, although achieving lower performance as expected due to lower parameter count and compute expenditure per environment step. The parameter count for the $\text{DRC}(1,1)$ and other networks used in the paper are reported in Table 2.
Figure 9 shows the performance of the $\text{DRC}(1,1)$ network on the medium difficulty validation levels during training. The $\text{DRC}(1,1)$ network shows a similar planning effect as the $\text{DRC}(3,3)$ network (Figure 1), which matches the results of Guez et al. [2019] who also showed that extra thinking steps improved performance of both the $\text{DRC}(1,1)$ and the $\text{DRC}(3,3)$ network.
We train the box-directions probe on the $\text{DRC}(1,1)$ network and evaluate it on the medium validation levels. The resulting probe achieves an F1 score of $72.3\%$. The F1 score is high enough to suggest that the $\text{DRC}(1,1)$ network also has a plan representation similar to the $\text{DRC}(3,3)$ network. Upon inspecting the probe visualizations, we find that the probe is indeed good at predicting the box movements in the short-term but is worse than the probe on $\text{DRC}(3,3)$ at predicting the long-term box movements (Figure 12, left). This suggests that the $\text{DRC}(1,1)$ network has a plan-representation, but it is limited to short-term planning. This results matches the evidence from the network’s overall performance which is significantly worse on level requiring long-term planning.
We now ask whether the performance improvement during thinking steps can be explained by the network refining its plan as we showed for the $\text{DRC}(3,3)$ network in Section 4. We measure the plan quality as we did in the main text using the length of plan from boxes and number of positive predictions. We find that the plan quality (Figure 12, middle) as well as the F1 score (Figure 12, right) measured using the probe increases over the number of thinking steps. Figure 11 (left) shows that this network also avoids myopic strategies when given extra thinking steps. This shows that the $\text{DRC}(1,1)$ network also refines its plan during the thinking steps.
Finally, we show that the $\text{DRC}(1,1)$ network also exhibits the pacing behavior as the $\text{DRC}(3,3)$ network. Figure 10 (left) shows that $\text{DRC}(1,1)$ also performs cycles at the start of levels, where extra computation steps are more useful to come up with better solutions. Figure 10 (right) shows that 60% of these cycles disappear with extra thinking steps, suggesting that they are performed by deliberately and not accidentally. The percentage is lower than 75% for $\text{DRC}(3,3)$ suggesting that $\text{DRC}(1,1)$ performs comparatively more accidental cycles, which is again consistent with lower planning-capacity and performance of the network.
Hence, the main Findings 1--3 analysed for $\text{DRC}(3,3)$ in the main text are also true for $\text{DRC}(1,1)$, with results suggesting worse long-term planning as compared to $\text{DRC}(3,3)$ as expected.
Figure 9: The $\text{DRC}(1,1)$ network shows similar planning effect as the $\text{DRC}(3,3)$. The plot is computed on the medium difficulty validation levels. The planning effect is measured as the difference in performance between the best thinking step and no thinking step. It achieves a lower performance than the $9$ layer ResNet network due to lower parameter count.
Figure 10: Left: $\text{DRC}(1,1)$ also performs cycles at the start of levels when it is more beneficial. Right: $\text{DRC}(1,1)$ performs many cycles when given no thinking steps. More than 60% of the cycles disappear with extra thinking steps.
Figure 11: Left: Thinking steps also help $\text{DRC}(1,1)$ to avoid myopic plans, thus increasing the time to place boxes first 3 boxes in solved levels while solving the level faster by placing the box 4 earlier. Right: The average optimal length of level solved by extra thinking steps doesn’t change with more thinking steps. This suggests that, unlike $\text{DRC}(3,3)$ which solves more difficult levels with more thinking steps, $\text{DRC}(1,1)$’s planning capacity is limited.
Figure 12: Left: For $\text{DRC}(1,1)$, the number of steps in advance that the Box-Directions probe is able to predict a box move shows similar trend as $\text{DRC}(3,3)$, but is overall lower, indicating that the $\text{DRC}(1,1)$ is less capable of planning. Middle and Right: The plan quality, as measured by summing probe-predicted chain lengths starting from boxes and counting squares with positive non-empty predictions, and the F1-score increases over thinking steps (Section 3.1). This suggests that $\text{DRC}(1,1)$ also refines its plan during computation, although it has an overall lesser F1 score of $60\%$ as compared to $84\%$ for $\text{DRC}(3,3)$.
Figure 13: Left: The box-directions probe trained on all layers for the ResNet model can predict moves with $80\%$ accuracy 50 steps before they occur. This implies that the ResNet model also computes a long-term plan for every observation. Middle: The plan quality and accuracy measured with the all-layer box-directions probe doesn’t improve with more environment steps as expected, since the plan in the activations need to be recomputed entirely from scratch on every step. This is unlike the DRC models which improve their plans with more thinking and environment steps. Right: Training the box-directions probe on each layer separately reveals that the plan representation is computed and improved in the early middle layers 1-4 with F1-score reaching 63.9%. The layer likely compute different parts of the plan since the probe trained on all the layers achieves a significantly higher F1-score of 84.5%.
B.2 ResNet
We now check if our findings about the DRC networks also generalize to a ResNet model with no recurrence over hidden state. The training and architecture details of the ResNet model is provided in Appendix A.
We first train the box-directions probe on the concatenation of the output of the ReLU activations from every residual block in every layer. This creates a single input vector of size $1024$ for the linear probe. The activations are collected on 1000 uniformly sampled levels from the medium-difficulty train set on which the ResNet model has a solve rate of $59.9\%$. The resulting probe achieves an F1 score of $84.5\%$ for predicting future box-movement directions on the validation set, similar to the probe on the $\text{DRC}(3,3)$ network. Figure 13 (left) also shows that the activations of the model in early steps are highly predictive of the actions it will take many steps in the future. This suggests that both the networks have similar representations that store their long-term plans. However, since the ResNet model achieves a lower performance on medium and hard difficulty sets, we can conclude that the ResNet model is not as good at constructing correct plans as the $\text{DRC}(3,3)$ network.
We found that plan improves with more thinking steps for the DRC model. Since the ResNet is not a recurrent model, it cannot take in extra thinking steps like the DRC networks can. Since the ResNet does not have a recurrent hidden state, we would expect that the plan representation of the model should not improve with more environments steps, as the activations are recomputed from scratch on every step. Figure 13 (middle plots) show that the plan is computed in one-step, and more environment steps do not result in better plan representations. Figure 13 (right) shows the validation F1-score of the same probe trained on the activations of the individual layers separately. We can see that the plan improves in the early to middle layers, implying that the plan representations are sequentially through the layers. Since the ResNet is state-less, it cannot exhibit the pacing-behavior.
Hence, we can conclude that only the finding 1 of plan representation holds for the ResNet model, with plan-improvements happening sequentially through the layers.
Appendix C Generalizing the $\text{DRC}(3,3)$ to larger levels
@@
We license the levels by Þorsteinsson [2009] as GPLv3, and make them available in the supplementary material. Table 4 shows the performance of DRC acriss various level sets. Higher scores align with levels considered easier by humans; for example, “Dimitri & Yorick” was made for children by Jacques Duthen, features small levels (maximum size: $12\times 10$) with at most 5 boxes. Level sets where the DRC fails to solve any puzzles were also challenging for the authors.
@@
We tested the $\text{DRC}(3,3)$ with 2 to 128 extra thinking steps, incrementing in powers of two. For most sets, we find some benefit to 2-4 thinking steps, but no more. The sole exception is XSokoban, which contains one level (Figure 15(c)) requiring 128 steps of thinking to solve.
@@
Table 4: Performance of the $\text{DRC}(3,3)$ on each set of levels by [Þorsteinsson, 2009]. The “max solved” columns represent the proportion of levels solved at the number of steps in the “max at” column, which is the highest solved proportion for each number of thinking steps tried. The “largest level size” column represents the height and width of the solved level with the largest grid area.
All levels Levels larger than $10\times 10$ Largest
Level collection # solved Max slv. max at # solved Max slv. max at level size
Dimitri & Yorick 61 86.9% 93.4% 2 0 (10, 12)
Sokoban Jr. 1 60 85.0% 85.0% 0 19 73.7% 73.7% 0 (18, 21)
Howard’s 3rd set 40 70.0% 70.0% 0 1 0.0% 0.0% 0 (11, 10)
Simple sokoban 61 55.7% 62.3% 16 51 47.1% 54.9% 16 (16, 19)
Sokoban Jr. 2 54 48.1% 53.7% 2 40 45.0% 47.5% 2 (14, 27)
Sokogen 990602 78 37.2% 43.6% 4 0 (10, 10)
Microban 155 31.0% 31.6% 64 17 5.9% 11.8% 2 (14, 10)
Yoshio Automatic 52 28.8% 36.5% 2 0 (10, 10)
Deluxe 55 25.5% 27.3% 16 1 0.0% 0.0% 0 (10, 13)
Howard’s 2nd set 40 12.5% 15.0% 2 22 0.0% 0.0% 0 (11, 10)
Sasquatch III 16 6.2% 6.2% 0 8 0.0% 0.0% 0 (10, 17)
Microcosmos 40 5.0% 10.0% 16 0 (10, 10)
Howard’s 1st set 100 4.0% 4.0% 0 54 0.0% 0.0% 0 (12, 10)
Still more levels 35 2.9% 2.9% 0 34 2.9% 2.9% 0 (13, 13)
Sasquatch IV 36 2.8% 2.8% 0 20 0.0% 0.0% 0 (10, 10)
Xsokoban 40 2.5% 5.0% 128 39 2.6% 5.1% 128 (13, 19)
Sasquatch 49 2.0% 4.1% 4 39 0.0% 2.6% 4 (14, 24)
@@
C.1 Generalization to zig-zag levels
Figure 16: $16\times 16$ zig-zag level that the $\text{DRC}(3,3)$ doesn’t solve. The network solves all levels of this pattern only up to $15\times 15$ size and fails on all zig-zag levels beyond that.
We create a template of levels as shown in Figure 16 where the agent has to take boxes from the left side to the targets on the right side by going through multiple zig-zag vertical lanes. This template can be created for arbitrary sized levels by adjusting the number of vertical lanes that the agent has to pass through. We find that the agent is able to solve levels only until the size of $15\times 15$ with 5 lanes but fails to solve larger levels. In larger levels, the agent does push the boxes through some of the lanes but is unable to find the complete path to the target through the remaining zig-zag lanes.
Appendix D Bistable and unstable plans in toy environments
Figure 17: Left: A custom level where the agent has two equally good paths to follow. The arrows show the prediction of the Agent-Directions probe with opacity proportional to the number of times an arrow was predicted across all the steps. The green and red arrows are correct and incorrect predictions, respectively. The agent behaves suboptimally by returning to the start and going left after first going right for three steps. Middle: The agent takes the path on the right after intervening with the corresponding arrows using the Agent-Directions probe on the first step. The same happens on the left if we intervene on that path without the suboptimal steps right of the undisturbed agent. Right: An empty level with Box-Directions probe intervening on the first four steps which are followed correctly by the agent on those four steps. When the intervention is removed on the fifth step, the agent computes the simpler path in green and doesn’t follow the path laid out earlier (in red).
We examined the Box-Directions probe, which displayed significantly higher causal influence than the Agent-Directions probe Table 1. To better understand this behavior, we designed controlled scenarios. Figure 17 (left) shows a level where a single box is next to a target, and the agent faces two equally viable paths. Without intervention, the agent initially chooses the right, taking three steps before returning to the start and switching to the left. This behavior deviates from optimality, which requires committing to a single path.
When we applied Agent-Directions probe to enforce a path choice, the agent consistently followed the chosen trajectory Figure 17 (middle). However, in levels where multiple viable paths exists, the Agent-Directions probe’s influence diminished, and the Box-Directions probe guided the agent more effectively. Figure 17 (right) shows a near-empty level where the Box-Directions probe was used to guide the agent on the first four steps. After the intervention ceased. the agent computed a shorter, optimal path and deviated from the initial intervention.
This analysis the agent prioritizes box-directions when available and relies on the agent-direction cues as secondary guidance. These findings highlight the hierarchy of planning components encoded in the network.
Appendix E Case Studies